# PuzzleFusion++: Auto-agglomerative 3D Fracture Assembly by Denoise and Verify

### v1

@@ Abstract
This paper proposes a novel “auto-agglomerative” 3D fracture assembly method, PuzzleFusion++, resembling how humans solve challenging spatial puzzles. Starting from individual fragments, the approach 1) aligns and merges fragments into larger groups akin to agglomerative clustering and 2) repeats the process iteratively in completing the assembly akin to auto-regressive methods. Concretely, a diffusion model denoises the 6-DoF alignment parameters of the fragments simultaneously, and a transformer model verifies and merges pairwise alignments into larger ones, whose process repeats iteratively. Extensive experiments on the Breaking Bad dataset show that PuzzleFusion++ outperforms all other state-of-the-art techniques by significant margins across all metrics In particular by over 10% in part accuracy and 50% in Chamfer distance. The code will be available on our project page: https://puzzlefusion-plusplus.github.io.
*Equal contribution.
@@ 1 Introduction
Deep neural networks (DNNs) have brought dramatic progress in developing such an intelligent computational method. DNN-based encoder enables robust shape matching by encoding raw 3D scans (i.e., point clouds) into latent embeddings, where a combinatorial optimization technique searches for an optimal set of pairwise alignments [18]. PuzzleFusion [13] proposed an innovative fully neural system based on Diffusion Models that simultaneously aligns all the pieces in solving 2D jigsaw puzzles. This paper pushes the frontier of spatial puzzle solving by proposing a fully neural system that simulates how humans tackle this problem.
Starting from individual fracture fragments, the approach 1) simultaneously aligns and merges fragments into larger groups, akin to agglomerative clustering, and 2) repeats this process iteratively in completing the assembly akin to auto-regressive methods. We call our approach “auto-agglomerative”, mimicking the way humans tackle challenging spatial puzzles. Concretely, after using PointNet++ [24] and VQVAE [37] to encode each fragment into latent embeddings, the approach uses a diffusion model to solve for the 6-DoF alignments of all the fragments. A transformer model verifies the inferred pairwise alignments and merges verified pairs into larger groups. The process repeats until forming a single assembly.
Extensive experiments on the Breaking Bad dataset [30] show that PuzzleFusion++ outperforms all existing methods by significant margins in all metrics, specifically by over 10% in part accuracy and over 50% in the Chamfer distance metric, pushing the frontier of computational methods for challenging spatial puzzle tasks.
@@ 2 Related Work
Diffusion models. Diffusion models [12, 29, 32] have gained prominence for their exceptional performance in generating images [26, 27, 9], videos [3, 4], and 3D assets [43, 19, 23, 34, 2]. The great capacity to capture complex data distribution makes diffusion-based models suitable for many tasks beyond generation. Recent works have extended diffusion models as general solvers for deterministic tasks, including object detection [6], camera pose estimation [38], shape reconstruction [8, 5], and puzzle solving [13, 28]. Our method exploits diffusion models within an agglomerative puzzle-solving framework to deal with the complex geometric reasoning for the fracture assembly tasks.
@@
Auto-regressive methods. Auto-regressive methods recursively predict the next values in a sequence based on the previously observed or predicted elements. Auto-regressive approaches like GPT [25] and its follow-ups have revolutionized natural language processing (NLP). Similar approaches also generate images as pixel sequences [36], audio [35], or 3D shapes [22, 31], etc. The auto-regressive agglomeration process in our method draws inspiration from these successes. By iteratively aligning and merging fractured fragments into larger ones for future iterations, we gradually assemble all fragments into a single object. The process mimics human cognitive puzzle-solving strategies, aiming for more precise and efficient assembly.
@@
3D shape assembly. 3D fracture assembly has been a challenge for machine intelligence [14]. PartNet [20] presents a large 3D object segmentation database (e.g., the legs or seat of a chair), fueling research on learning-based methods [42, 17, 21]. Recent transformer-based approaches [41, 11, 45] capture the global semantic and geometric contexts well to enhance the alignment accuracy. Chen et al. [7] proposes a shape-mating dataset where objects are randomly fractured into two fragments, and an adversarial learning framework reassembles the two fragments. Lamb et al. [15] collects a dataset with real-world broken objects. The Breaking Bad dataset [30] introduces a more challenging benchmark, where diverse objects are fractured into multiple (i.e., 2 to 100) fragments via physical simulations. Jigsaw [18] utilizes deep networks for fracture surface segmentation and matching, followed by a global alignment step to derive the poses. SE(3)-equiv [39] extracts global equivariant and invariant features to represent each fractured fragment and directly regress the 6-DoF alignment parameters. DiffAssemble [28] further applies a diffusion model to this framework. This paper introduces an auto-agglomerative framework to mimic how humans solve spatial puzzles, where a diffusion model with local geometric features serves as a spatial solver in each iteration, where a transformer verifies the correctness of aligned fragments and merges them into larger groups.
@@ 3 PuzzleFusion++: An auto-agglomerative approach for fracture assembly
Figure 2: The architecture overview of PuzzleFusion++. Left: An illustration of the auto-agglomerative fracture assembly process with the first two iterations. Right: Close-ups of the SE3 denoise transformer and the pairwise alignment verifier transformer. Please refer to Figure 3 for the details of the architectures.
@@ 3.1 Preliminaries
Input representation. An input fragment comes as point clouds. We train PointNet++ [24] and VQ-VAE [37] to encode the point clouds into latent vectors at the points (details in Section A.1). PointNet++ employs furthest point sampling to select 25 points, creating 25 corresponding point latent vectors for each fragment. This approach retains intricate surface details, offering an advantage over the global feature extraction used in previous studies [39, 28]. Note that the encoding process normalizes the 3D coordinate frame so that the center of mass is at the origin, and the longest dimension of the axis-aligned bounding box becomes a unit length. The encoding is still sensitive to rotation, where latent vectors are re-calculated as we optimize fragment alignments.
@@
Anchor fragments. Fracture assembly has an inherent ambiguity in the 3D rigid transformation. We introduce the concept of anchor fragments, whose alignments are fixed to the identity matrix for rotation and the zero vector for translation. During inference, the largest fragment (based on the longest dimension of its axis-aligned bounding box) becomes an anchor. Any fragment that merges with an anchor also becomes an anchor. During training, we set anchors to be the largest fragment (with the same definition as above) plus its neighboring fragments with a 50% chance each to simulate the merging process at the inference time.
@@ 3.2 SE3 denoiser
Figure 3: Architecture specifications. The denoiser (blue) is a diffusion model with Transformer architecture at its core. The verifier (orange) is a Transformer.
Forward process. Let $x^{f}_{t}$ denote the 7D alignment vector of fragment $f$ at timestep $t$. $x^{f}_{0}$ is the ground-truth alignment before noise injection. A piece-wise quadratic scheduler adds a Gaussian Noise $\delta^{f}_{t}$ to $x^{f}_{0}$, using $m$=700 and $T$=1000 (refer to Figure 3). No noise is added to anchor fragments.
Reverse process. Following PuzzleFusion [13], each sampled point of each fragment keeps track of the alignment estimate in the denoise transformer architecture. The denoising architecture is standard (See Figures 2 and 3), where the main components are 1) Intra-fragment self-attention (Local SA) among 25 sampled points within a fragment; 2) Inter-fragment self-attention (Global SA) among all sampled points across all fragments; and 3) Average pooling over the 25 sampled points of fragment $f$ to predict the residual $\tilde{\delta}^{f}_{t}$.
Both intra-fragment and inter-fragment SA modules utilize adaptive layer norm [10] to inject the timestep $t$. Each module comprises 6 self-attention layers, each with 8 multi-heads and a feature dimension of 512.
@@
Note that anchor fragments are processed exactly the same way, while their alignment should be the identity rotation and zero-vector translation. Therefore, alignment estimation is discarded at every denoising step during inference, and no gradients are injected during training for anchor fragments. The standard DDPM [12] loss is imposed on non-anchor fragments: $E_{t,x_{0}^{f},\delta_{t}^{f}}\left\lVert\delta^{f}_{t}-\tilde{\delta}^{f}_{t}\right\rVert^{2}$.
@@ 3.3 Pairwise alignment verifier
An input node embedding for a pair of fragments ($f$ and $g$) is the sum of two vectors. The first vector is the concatenation of the sinusoidal positional encodings of the fragment indices, $[PE(f),PE(g)]\in\mathbb{R}^{256}$. The second vector represents the alignment quality of $f$ and $g$, utilizing a point matching module from Jigsaw by Lu et al. [18]. Specifically, the module takes the point clouds of all the fragments (without alignment) and identifies a set of matching points. For point matches between $f$ and $g$, we calculate the Euclidean distances of points and construct a normalized histogram with six bins. We append the total number of matches as the seventh value, which is fed into an MLP to construct the second vector. The distance thresholds of these bins are set at (0, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1, $\infty$). The verifier is optimized using the standard binary cross-entropy loss.
@@ 3.4 Auto-agglomerative inference
The denoiser and the verifier iteratively align and merge fracture fragments into larger groups. This process repeats six iterations or until all the fragments are merged into a single component. Fragment pairs are verified for merging if their classification score exceeds a threshold of 0.9. One challenge in fragment merging is the potential inclusion of points from inner surfaces when taking the union of the point-clouds. We employ simple heuristics to detect and remove such inner points before merging. Specifically, we compute a surface normal for each point of a point-cloud using estimate_pointcloud_normals in the pytorch3d.ops module. A point is identified as an inner surface point if another point in the opposing point-cloud is within a distance of 0.001 and their surface normals yield a negative dot product. A point-cloud of a fragment contains 1,000 points. After removing inner surface points, we resample to maintain 1,000 points using the farthest point sampling method (§4). For anchor fragments, we preserve them as separate and freeze their alignments rather than merging to retain the high-resolution information crucial for large (e.g., anchor) fragments.
@@ 4 Experiments
Datasets. Following recent works [39, 18], we use the Breaking Bad dataset [30]. Specifically, 34,075 assemblies from 407 objects in the everyday subset are for training. 7,679 assemblies from 91 objects in the everyday subset and 3,651 assemblies from 40 uncategorized objects in the artifact subset are for testing. We limit assemblies with at most 20 fragments and do not use the assembly category labels (e.g., bottles, vases, and mirrors). Note that the training/testing split is exactly the same as Breaking Bad dataset [30] for fair comparisons.
Evaluation metrics. The Breaking Bad dataset offers three evaluation metrics, where the anchor fragment aligns the predicted assembly and the ground truth before the metric calculation:
@@
$\bullet$ Global, LSTM, and DGL are the initial baselines provided by the Breaking Bad dataset. Global combines the global shape features with per-fragment features and directly regresses the pose. LSTM focuses on learning cross-fragment relationships, using a bi-directional LSTM to predict pose. DGL utilizes the graph neural network to find the relationship between fragments. These methods are originally designed for PartNet [20] assembly tasks.
$\bullet$ SE(3)-Equiv [39] integrates equivariant and invariant features to model multi-part correlations. The released code only trains and tests on samples with at most 8 fragments. We train and test their system for assemblies with at most 20 fragments for fair comparison.
$\bullet$ DiffAssemble [28] uses the equivariant encoder to extract per-fragment features similar to SE(3) and uses a diffusion model to predict the pose.
$\bullet$ Jigsaw [18] is the state-of-the-art fracture assembly method, leveraging hierarchical features of global and local geometry to do fracture surface point matching and recover the poses.
[…]
Trained and tested on the everyday subset
Global[16, 29] 80.7 15.1 24.6 14.6 23.7
LSTM[40] 84.2 16.2 22.7 15.8 29.9
DGL[44] 79.4 15.0 31.0 14.3 28.7
SE(3)-Equiv[39] 79.3 16.9 8.41 28.5 129.9
DiffAssemble[28] 73.3 14.8 27.5 - -
Jigsaw[18] 42.3 10.7 57.3 13.3 1063.5
PuzzleFusion++ (Ours) 38.1 8.04 70.6 6.03 928.5
Trained on the everyday subset, tested on the artifact subset
Jigsaw[18] 52.4 22.2 45.6 14.3 -
PuzzleFusion++ (Ours) 52.1 13.9 49.6 14.5 -

### v2

@@ Abstract
This paper proposes a novel “auto-agglomerative” 3D fracture assembly method, PuzzleFusion++, resembling how humans solve challenging spatial puzzles. Starting from individual fragments, the approach 1) aligns and merges fragments into larger groups akin to agglomerative clustering and 2) repeats the process iteratively in completing the assembly akin to auto-regressive methods. Concretely, a diffusion model denoises the 6-DoF alignment parameters of the fragments simultaneously, and a transformer model verifies and merges pairwise alignments into larger ones, whose process repeats iteratively. Extensive experiments on the Breaking Bad dataset show that PuzzleFusion++ outperforms all other state-of-the-art techniques by significant margins across all metrics In particular by over 10% in part accuracy and 50% in Chamfer distance. The code and models are available on our project page: https://puzzlefusion-plusplus.github.io.
@@ 1 Introduction
Deep neural networks (DNNs) have brought dramatic progress in developing such an intelligent computational method. DNN-based encoder enables robust shape matching by encoding raw 3D scans (i.e., point clouds) into latent embeddings, where a combinatorial optimization technique searches for an optimal set of pairwise alignments (Lu et al., 2023). PuzzleFusion (Hossieni et al., 2023) proposed an innovative fully neural system based on Diffusion Models that simultaneously aligns all the pieces in solving 2D jigsaw puzzles. This paper pushes the frontier of spatial puzzle solving by proposing a fully neural system that simulates how humans tackle this problem.
Starting from individual fracture fragments, the approach 1) simultaneously aligns and merges fragments into larger groups, akin to agglomerative clustering, and 2) repeats this process iteratively in completing the assembly akin to auto-regressive methods. We call our approach “auto-agglomerative”, mimicking the way humans tackle challenging spatial puzzles. Concretely, after using PointNet++ (Qi et al., 2017) and VQVAE (Van Den Oord et al., 2017) to encode each fragment into latent embeddings, the approach uses a diffusion model to solve for the 6-DoF alignments of all the fragments. A transformer model verifies the inferred pairwise alignments and merges verified pairs into larger groups. The process repeats until forming a single assembly.
Extensive experiments on the Breaking Bad dataset (Sellán et al., 2022) show that PuzzleFusion++ outperforms all existing methods by significant margins in all metrics, specifically by over 10% in part accuracy and over 50% in the Chamfer distance metric, pushing the frontier of computational methods for challenging spatial puzzle tasks. The code and models are available on our project page: https://puzzlefusion-plusplus.github.io.
@@ 2 Related Work
Diffusion models. Diffusion models (Ho et al., 2020; Song et al., 2020; Sohl-Dickstein et al., 2015) have gained prominence for their exceptional performance in generating images (Ramesh et al., 2022; Rombach et al., 2022; Dhariwal & Nichol, 2021b), videos (Blattmann et al., 2023a; b), and 3D assets (Zeng et al., 2022; Luo & Hu, 2021; Poole et al., 2022; Tang et al., 2024; Alliegro et al., 2023). The great capacity to capture complex data distribution makes diffusion-based models suitable for many tasks beyond generation. Recent works have extended diffusion models as general solvers for deterministic tasks, including object detection (Chen et al., 2023b), camera pose estimation (Wang et al., 2023), shape reconstruction (Cheng et al., 2023; Chen et al., 2023a), and puzzle solving (Hossieni et al., 2023; Scarpellini et al., 2024). Our method exploits diffusion models within an agglomerative puzzle-solving framework to deal with the complex geometric reasoning for the fracture assembly tasks.
@@
Auto-regressive methods. Auto-regressive methods recursively predict the next values in a sequence based on the previously observed or predicted elements. Auto-regressive approaches like GPT (Radford et al., 2018) and its follow-ups have revolutionized natural language processing (NLP). Similar approaches also generate images as pixel sequences (Van Den Oord et al., 2016b), audio (Van Den Oord et al., 2016a), or 3D shapes (Nash et al., 2020; Siddiqui et al., 2023), etc. The auto-regressive agglomeration process in our method draws inspiration from these successes. By iteratively aligning and merging fractured fragments into larger ones for future iterations, we gradually assemble all fragments into a single object. The process mimics human cognitive puzzle-solving strategies, aiming for more precise and efficient assembly.
@@
3D shape assembly. 3D fracture assembly has been a challenge for machine intelligence (Huang et al., 2006). PartNet (Mo et al., 2019) presents a large 3D object segmentation database (e.g., the legs or seat of a chair), fueling research on learning-based methods (Yin et al., 2020; Li et al., 2023; Narayan et al., 2022). Recent transformer-based approaches (Xu et al., 2024; Du et al., 2024; Zhang et al., 2022) capture the global semantic and geometric contexts well to enhance the alignment accuracy. Chen et al. (2022) proposes a shape-mating dataset where objects are randomly fractured into two fragments, and an adversarial learning framework reassembles the two fragments. Lamb et al. (2023) collects a dataset with real-world broken objects. The Breaking Bad dataset (Sellán et al., 2022) introduces a more challenging benchmark, where diverse objects are fractured into multiple (i.e., 2 to 100) fragments via physical simulations. Jigsaw (Lu et al., 2023) utilizes deep networks for fracture surface segmentation and matching, followed by a global alignment step to derive the poses. SE(3)-equiv (Wu et al., 2023) extracts global equivariant and invariant features to represent each fractured fragment and directly regress the 6-DoF alignment parameters. DiffAssemble (Scarpellini et al., 2024) further applies a diffusion model to this framework. This paper introduces an auto-agglomerative framework to mimic how humans solve spatial puzzles, where a diffusion model with local geometric features serves as a spatial solver in each iteration, and a transformer verifies the correctness of aligned fragments and merges them into larger groups.
@@ 3 PuzzleFusion++: An auto-agglomerative approach for fracture assembly
Figure 2: The architecture overview of PuzzleFusion++ (mesh used only for visualization). Left: An illustration of the auto-agglomerative fracture assembly process with the first two iterations. Right: Close-ups of the SE3 denoise transformer and the pairwise alignment verifier transformer. Please refer to Figure 3 for the details of the architectures.
@@ 3.1 Preliminaries
Input representation. An input fragment comes as point clouds. We train PointNet++ (Qi et al., 2017) and VQ-VAE (Van Den Oord et al., 2017) to encode the point clouds into latent vectors at the points (details in Section A.1). PointNet++ employs furthest point sampling to select 25 points, creating 25 corresponding point latent vectors for each fragment. This approach retains intricate surface details, offering an advantage over the global feature extraction used in previous studies (Wu et al., 2023; Scarpellini et al., 2024). Note that the encoding process normalizes the 3D coordinate frame so that the center of mass is at the origin, and the longest dimension of the axis-aligned bounding box becomes a unit length. The encoding is still sensitive to rotation, where latent vectors are re-calculated as we optimize fragment alignments.
@@
Anchor fragments. Fracture assembly has an inherent ambiguity in the 3D rigid transformation. We introduce the concept of anchor fragments to settle down a reference coordinate system. The anchor fragments have fixed alignments (i.e., the identity matrix for rotation and the zero vector for translation) and do not move in the assembly process. At the beginning of the inference, the largest fragment (based on the longest dimension of its axis-aligned bounding box) becomes an anchor. In the assembly process, any fragment that merges with an anchor also becomes an anchor. During training, we set anchors to be the largest fragment (with the same definition as above) plus its neighboring fragments with a 50% chance each to simulate the inference-time merging process.
@@ 3.2 SE3 denoiser
Figure 3: Inference pipeline with architecture specifications. The denoiser (blue) is a diffusion model with Transformer architecture at its core. The verifier (orange) is a Transformer.
Forward process. Let $x^{f}_{t}$ denote the 7D alignment vector of fragment $f$ at timestep $t$. $x^{f}_{0}$ is the ground-truth alignment before noise injection. A piece-wise quadratic scheduler adds a Gaussian noise (except to the anchor fragment) $\delta^{f}_{t}$ to $x^{f}_{0}$, using $m$=700 and $T$=1000:
$x^{f}_{t}=\sqrt{\bar{\alpha}_{t}}x^{f}_{0}+\sqrt{1-\bar{\alpha}_{t}}\delta^{f}_{t},\quad\bar{\alpha}=\begin{cases}1-(0.1)\left(\frac{t}{m}\right)^{2}&\text{if }t\leq m\\ 0.9\left(1-\left(\frac{t-m}{T-m}\right)^{2}\right)&\text{if }m<t\leq T\end{cases}$ (1)
Reverse process. Following PuzzleFusion (Hossieni et al., 2023), each sampled point of each fragment keeps track of the alignment estimate in the denoise transformer architecture. The denoising architecture is standard (See Figures 2 and 3), where the main components are 1) Intra-fragment self-attention (Local SA) among 25 sampled points within a fragment; 2) Inter-fragment self-attention (Global SA) among all sampled points across all fragments; and 3) Average pooling over the 25 sampled points of fragment $f$ to predict the residual $\tilde{\delta}^{f}_{t}$.
Both intra-fragment and inter-fragment SA modules utilize adaptive layer norm (Dhariwal & Nichol, 2021a) to inject the timestep $t$. Each module comprises 6 self-attention layers, each with 8 multi-heads and a feature dimension of 512.
@@
Note that anchor fragments are processed exactly the same way, while their alignment should be the identity rotation and zero-vector translation. Therefore, alignment estimation is discarded at every denoising step during inference, and no gradients are injected during training for anchor fragments. The standard DDPM (Ho et al., 2020) loss is imposed on non-anchor fragments: $E_{t,x_{0}^{f},\delta_{t}^{f}}\left\lVert\delta^{f}_{t}-\tilde{\delta}^{f}_{t}\right\rVert^{2}$.
DiffAssemble (Scarpellini et al., 2024) applies diffusion models to solve both 2D and 3D reassembly problems but obtains suboptimal performance in the complex 3D fracture assembly task. Our denoiser provides tailored designs over all critical components of diffusion models. Please refer to §4.3 for detailed analyses and experimental results.
@@ 3.3 Pairwise alignment verifier
An input node embedding for a pair of fragments ($f$ and $g$) is the sum of two vectors. The first vector is the concatenation of the sinusoidal positional encodings of the fragment indices, $[PE(f),PE(g)]\in\mathbb{R}^{256}$. The second vector represents the alignment quality of $f$ and $g$, utilizing a point matching module from Jigsaw by Lu et al. (2023). Specifically, the module takes the point clouds of all the fragments (without alignment) and identifies a set of matching points. For point matches between $f$ and $g$, we calculate the Euclidean distances of points and construct a normalized histogram with six bins. We append the total number of matches as the seventh value, which is fed into an MLP to construct the second vector. The distance thresholds of these bins are set at (0, 1e-3, 5e-3, 1e-2, 5e-2, 1e-1, $\infty$). The verifier is optimized using the standard binary cross-entropy loss.
@@ 3.4 Auto-agglomerative inference
The denoiser and the verifier iteratively align and merge fracture fragments into larger groups. This process repeats six iterations or until all the fragments are merged into a single component. Fragment pairs are verified for merging if their classification score exceeds a threshold of 0.9. One challenge in fragment merging is the potential inclusion of points from inner surfaces when taking the union of the point-clouds. We employ simple heuristics to detect and remove such inner points before merging. Specifically, we compute a surface normal for each point of a point-cloud using estimate_pointcloud_normals in the pytorch3d.ops module. A point is identified as an inner surface point if another point in the opposing point-cloud is within a distance of 0.001 and their surface normals yield a negative dot product. A point-cloud of a fragment contains 1,000 points. After removing inner surface points, we resample to maintain 1,000 points using the farthest point sampling method. For anchor fragments, we preserve them as separate and freeze their alignments rather than merging to retain the high-resolution information crucial for large (e.g., anchor) fragments.
@@ 4 Experiments
Datasets. Following recent works (Wu et al., 2023; Lu et al., 2023), we use the Breaking Bad dataset (Sellán et al., 2022). Specifically, 34,075 assemblies from 407 objects in the everyday subset are for training. 7,679 assemblies from 91 objects in the everyday subset and 3,651 assemblies from 40 uncategorized objects in the artifact subset are for testing.
[…]
Trained and tested on the everyday subset
Global (Li et al., 2020) 80.7 15.1 24.6 14.6 23.7
LSTM (Wu et al., 2020) 84.2 16.2 22.7 15.8 29.9
DGL (Zhan et al., 2020) 79.4 15.0 31.0 14.3 28.7
SE(3)-Equiv(Wu et al., 2023) 79.3 16.9 8.41 28.5 129.9
DiffAsb(Scarpellini et al., 2024) 73.3 14.8 27.5 - -
Jigsaw(Lu et al., 2023) 42.3 10.7 57.3 13.3 1063.5
PuzzleFusion++ (Ours) 38.1 8.04 70.6 6.03 928.5
Trained on the everyday subset, tested on the artifact subset
Jigsaw(Lu et al., 2023) 52.4 22.2 45.6 14.3 -
PuzzleFusion++ (Ours) 52.1 13.9 49.6 14.5 -
We limit assemblies with at most 20 fragments and do not use the assembly category labels (e.g., bottles, vases, and mirrors). Each fragment is represented as a point cloud consisting of 1,000 points, following the baseline settings. Note that the training/testing split is exactly the same as Breaking Bad dataset (Sellán et al., 2022) for fair comparisons.
Evaluation metrics. The Breaking Bad dataset offers three evaluation metrics, where the biggest fragment is used to align the predicted assembly with the ground truth before calculating the metrics, which is the same as in Jigsaw (Lu et al., 2023) evaluation settings:
@@
$\bullet$ Global, LSTM, and DGL are the initial baselines provided by the Breaking Bad dataset. Global combines the global shape features with per-fragment features and directly regresses the pose. LSTM focuses on learning cross-fragment relationships, using a bi-directional LSTM to predict pose. DGL utilizes the graph neural network to find the relationship between fragments. These methods are originally designed for PartNet (Mo et al., 2019) assembly tasks.
$\bullet$ SE(3)-Equiv (Wu et al., 2023) integrates equivariant and invariant features to model multi-part correlations. The released code only trains and tests on samples with at most 8 fragments. We train and test their system for assemblies with at most 20 fragments for fair comparison.
$\bullet$ DiffAssemble (Scarpellini et al., 2024) uses the equivariant encoder to extract per-fragment features similar to SE(3) (Wu et al., 2023) and uses a diffusion model to predict the pose.
$\bullet$ Jigsaw (Lu et al., 2023) is the state-of-the-art fracture assembly method, leveraging hierarchical features of global and local geometry to do fracture surface point matching and recover the poses.